# Rendering Wireless Environments Useful for Gradient Estimators: A Zero-Order Stochastic Federated Learning Method

### v1

@@ III Convergence analysis
We begin by assuming that a global minimizer $\theta^{*}\in\mathbb{R}^{d}$ exists such that $\min_{\theta\in\mathbb{R}^{d}}F(\theta)=F(\theta^{*})>-\infty$ and $\nabla F(\theta^{*})=0$.
@@
We assume the existence and the continuity of $\nabla F_{i}(\theta)$ and $\nabla^{2}F_{i}(\theta)$, and that there exists a constant $\alpha_{1}>0$ such that $\|\nabla^{2}F_{i}(\theta)\|_{2}\leq\alpha_{1}$,$\forall i\in\mathcal{N}$.
@@
By Assumption 4, we know that the objective function $\theta\longmapsto F(\theta)$ is $L$-smooth for some positive constant $L$, $\|\nabla F(\theta)-\nabla F(\theta^{\prime})\|\leq L\|\theta-\theta^{\prime}\|,\;\forall\theta,\theta^{\prime}\in\mathbb{R}^{d},$ or equivalently, $F(\theta)\leq F(\theta^{\prime})+\langle\nabla F(\theta^{\prime}),\theta-\theta^{\prime}\rangle+\frac{L}{2}\|\theta-\theta^{\prime}\|^{2}$.
@@
By Assumptions 1, 2, and 4, we can find a scalar value $c_{3}>0$ such that $\|b_{k}\|\leq c_{3}\gamma_{k}$. Proof: Refer to Appendix B-C.
@@
To satisfy Assumption 5, we consider the following form of the step sizes, $\alpha_{k}=\alpha_{0}(1+k)^{-\upsilon_{1}}$ and $\gamma_{k}=\gamma_{0}(1+k)^{-\upsilon_{2}}$ with $\upsilon_{1},\upsilon_{2}>0$. Then, it is sufficient to find $\upsilon_{1}$ and $\upsilon_{2}$ such that $0<\upsilon_{1}+\upsilon_{2}\leq 1$, $\upsilon_{1}+3\upsilon_{2}>1$, and $\upsilon_{1}>0.5$.
@@
If all Assumptions 1-5 hold and $\|\theta_{k}\|<\infty$ almost surely, then for any constant $\nu>0$, we have $\mathbb{P}(\lim_{K\rightarrow\infty}\sup_{K^{\prime}\geq K}\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu)=0$. Proof: Refer to Appendix C-A.
@@
When Assumptions 1-5 hold, we have $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}<+\infty$ and $\lim_{k\rightarrow\infty}\|\nabla F(\theta_{k})\|=0$ almost surely, meaning that the algorithm converges. Proof: Refer to Appendix C-B.
Proof sketch: We substitute the algorithm’s updates in the second inequality of Lemma 3 and replace the estimate by its expectation plus the stochastic error. We then perform a recursive addition over the iterations $k>0$. With Lemma 5, the conditions on the step sizes, and the upper bound estimate’s squared norm, we are able to find an upper bound on $\sum_{k}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}$ when $k$ grows to $\infty$. The next step is to consider the hypothesis $\lim_{k\rightarrow\infty}\sup\|\nabla F(\theta_{k})\|\geq\rho$, for $\rho>0$, and prove that it contradicts with the first result.
Define $\delta_{k}=F(\theta_{k})-F(\theta^{*})$. We next find an upper bound on the convergence rate of Algorithm 1.
@@
Consider in addition to the assumptions in Theorem 1, that the step sizes are those of Example 2 with $\upsilon_{3}=\upsilon_{1}+\upsilon_{2}<1$. Then, we can write
$\begin{split}\frac{\sum_{k}\alpha_{k}\gamma_{k}\mathbb{E}\big{[}\|\nabla F(\theta_{k})\|^{2}\big{]}}{\sum_{k}\alpha_{k}\gamma_{k}}\leq\frac{A(1-\upsilon_{3})}{(K+2)^{1-\upsilon_{3}}-1}.\\ \end{split}$ (4)
with $A=\frac{2\delta_{0}}{c_{1}\alpha_{0}\gamma_{0}}+c_{3}^{2}\gamma_{0}^{2}\big{(}\frac{\upsilon_{1}+3\upsilon_{2}}{\upsilon_{1}+3\upsilon_{2}-1}\big{)}+\frac{Lc_{2}\alpha_{0}}{c_{1}\gamma_{0}}\big{(}\frac{2\upsilon_{1}}{2\upsilon_{1}-1}\big{)}$. Proof: Refer to Appendix C-C.
In Theorem 2, we see that the optimal choice of the exponents for the time-varying component $O(\frac{1}{K^{1-\upsilon_{1}-\upsilon_{2}}})$, is $\upsilon_{1}=\frac{1}{2}$ and $\upsilon_{2}=\frac{1}{6}$ for a rate of $O(\frac{1}{\sqrt[3]{K}})$. However, to prevent the constant component from growing too large, it is recommended to choose slightly larger exponents of $\upsilon_{1}=\frac{1}{2}+\frac{\epsilon}{2}$ and $\upsilon_{2}=\frac{1}{6}+\frac{\epsilon}{2}$, where $\epsilon$ is a small strictly positive value. This will result in a rate of $O\big{(}\frac{1}{K^{\frac{1}{3}-\epsilon}}\big{)}$.
III-B Non-Symmetrical Channels Case
Assuming a non-symmetrical channel model with $\mathbb{E}[h_{i,k}]=\mu_{h}$ and $\sigma_{h}^{2}=\mathbb{E}[h_{i,k}^{2}]-\mu_{h}^{2}$, $\forall i,\forall k$, we provide how our gradient estimates and algorithms can be adjusted in Appendix D to account for this case. In fact, non-symmetrical channel models (e.g., Rician) offer a simplification of both analysis and implementation in comparison to symmetrical models (e.g., Rayleigh), as the non-zero mean no longer cancels out the gradient, and the design is further independent of the autocorrelation of the channels.
@@ IV Experimental results
We classify photos of the two digits “00” and “$1$” from the MNIST dataset [37] using a nonconvex logistic regression model with a regularization parameter of $0.001$. All images are divided equally among the devices and are considered to be preprocessed by being compressed to have dimension $d=10$ using a lossy autoencoder. We run our code on $50$ simulations with different random model initializations testing the accuracy in every iteration against an independent test set. The graphs in Fig. 2 are averaged over all these simulations. For the non-IID data distribution, we first sort the images according to their labels and then divide them among the devices. $\Phi_{k}$ is generated according to Example 1. All channels are generated using the normal distribution with autocovariance $K_{hh}=\frac{1}{2}$. The noise is Gaussian with 00 mean and variance $\sigma_{n}^{2}=\frac{1}{4}$. The step size for FedAvg is taken as $\eta=0.15$. For 1P-ZOFL, $\alpha_{k}=0.5(1+k)^{-0.51}$ and $\gamma_{k}=2.5(1+k)^{-0.18}$. Our algorithm performs consistently well with all the different random variations influencing every simulation in line with our theoretical results. Considering non-IID data distribution seems to be better impacted by our ZO algorithm slightly at the beginning without a major effect on the final result.
In Fig. 3, we test our algorithm with different noise variance $\sigma_{n}^{2}$ and notice that while the algorithm is slowed slightly at the beginning, it does not affect the convergence. When $\sigma_{n}^{2}=2.25$, we change $\alpha_{0}=0.1$ and $\gamma_{0}=0.8$. When $\sigma_{n}^{2}=10.0489$, $\alpha_{0}=0.07$ and $\gamma_{0}=0.3$. We provide all other details alongside a discussion of our algorithm’s performance in relation to the noise variance in Appendix E.
@@ Appendix A On the channel model
$\begin{split}&z\sqrt{2}\cos(2\pi f_{c}t)\\ =&2x\sum_{i}a_{i}\cos(\varphi_{i}(t))\cos^{2}(2\pi f_{c}t)\\ &-2x\sum_{i}a_{i}\sin(\varphi_{i}(t))\sin(2\pi f_{c}t)\cos(2\pi f_{c}t)\\ &+\sqrt{2}w(t)\cos(2\pi f_{c}t)\end{split}$
$\begin{split}=&x\sum_{i}a_{i}\cos(\varphi_{i}(t))(1+\cos(4\pi f_{c}t))\\ &-x\sum_{i}a_{i}\sin(\varphi_{i}(t))\sin(4\pi f_{c}t)+\sqrt{2}w(t)\cos(2\pi f_{c}t)\\ \end{split}$ (7)
@@ B-A Proof of Lemma 1: Biased Estimator
$\begin{split}&\mathbb{E}[g_{k}|\mathcal{H}_{k}]\\ =&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}\Big{(}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(a)}{=}&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}h_{i,k+1}\tilde{F}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}\big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(b)}{=}&\mathbb{E}\Big{[}\Phi_{k}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{F}_{i}(\theta_{k})\\ &+\gamma_{k}\sum_{i=1}^{N}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}\Phi_{k}^{T}\nabla\tilde{F}_{i}(\theta_{k})\\ &+\gamma_{k}^{2}\sum_{i=1}^{N}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(c)}{=}&\mathbb{E}\Big{[}\Phi_{k}\bigg{(}\frac{\gamma_{k}}{\sigma_{h}^{2}}\sum_{i=1}^{N}h_{i,k+1}h_{i,k}\Phi_{k}^{T}\nabla\tilde{F}_{i}(\theta_{k})+\\ &+\gamma_{k}^{2}\sum_{i=1}^{N}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&\frac{\gamma_{k}}{\sigma_{h}^{2}}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}h_{i,k}\Big{|}\mathcal{H}_{k}\Big{]}\mathbb{E}\Big{[}\Phi_{k}\Phi_{k}^{T}\Big{|}\mathcal{H}_{k}\Big{]}\nabla\tilde{F}_{i}(\theta_{k})+\\ &\gamma_{k}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}\end{split}$
=$\begin{split}\overset{(d)}{=}&\gamma_{k}\alpha_{2}\frac{K_{hh}}{\sigma_{h}^{4}}\sum_{i=1}^{N}\nabla F_{i}(\theta_{k})+\\ &\gamma_{k}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(e)}{=}&c_{1}\gamma_{k}(\nabla F(\theta_{k})+b_{k}),\end{split}$ (9)
where $(a)$ is by the definition in (2) and due to Assumption 1, $(b)$ is by Taylor expansion and mean-value theorem and considering $\breve{\theta}_{k}$ between $\theta_{k}$ and $\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}$. $(c)$ is since $\mathbb{E}[h_{i,k+1}]=0$ for the first element and $\mathbb{E}[h_{i,k+1}h_{j,k}]=0$ when $i\neq j$ and the independence of noise for the second element. $(d)$ is due to Assumption 2. In $(e)$, we let $c_{1}=\alpha_{2}\frac{K_{hh}}{\sigma_{h}^{4}}$ and the bias $b_{k}=\gamma_{k}\frac{\sigma_{h}^{2}}{\alpha_{2}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}$.

### v2

@@ III Convergence analysis
We begin by assuming that a global minimizer $\theta^{*}\in\mathbb{R}^{d}$ exists such that $\min_{\theta\in\mathbb{R}^{d}}F(\theta)=F(\theta^{*})$ and $\nabla F(\theta^{*})=0$. We then define $\delta_{k}=F(\theta_{k})-F(\theta^{*})$ at every timeslot $k$.
@@
We assume the existence and the continuity of $\nabla F_{i}(\theta)$ and $\nabla^{2}F_{i}(\theta)$, and that there exists a constant $\beta_{3}>0$ such that $\|\nabla^{2}F_{i}(\theta)\|_{2}\leq\beta_{3}$,$\forall i\in\mathcal{N}$.
@@
By Assumption 4, we know that the objective function $\theta\longmapsto F(\theta)$ is $L$-smooth for some positive constant $L$, $\|\nabla F(\theta)-\nabla F(\theta^{\prime})\|\leq L\|\theta-\theta^{\prime}\|,\;\forall\theta,\theta^{\prime}\in\mathbb{R}^{d}$.
@@
By Assumptions 1, 2, and 4, we can find a scalar value $c_{3}>0$ such that $\|b_{k}\|\leq c_{3}\gamma_{k}$.
Proof: Refer to Appendix B-C.
@@
Consider $\alpha_{k}=\alpha_{0}(1+k)^{-\upsilon_{1}}$ and $\gamma_{k}=\gamma_{0}(1+k)^{-\upsilon_{2}}$ with $\upsilon_{1},\upsilon_{2}>0$. Then, it is sufficient to find $\upsilon_{1}$ and $\upsilon_{2}$ such that $0<\upsilon_{1}+\upsilon_{2}\leq 1$, $\upsilon_{1}+3\upsilon_{2}>1$, and $\upsilon_{1}>0.5$.
@@
If all Assumptions 1-5 hold, then for any constant $\nu>0$, we have $\mathbb{P}(\lim_{K\rightarrow\infty}\sup_{K^{\prime}\geq K}\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu)=0$.
Proof: Refer to Appendix C-A.
@@
When Assumptions 1-5 hold, then $\lim_{k\rightarrow\infty}\|\nabla F(\theta_{k})\|=0$ almost surely, meaning that the algorithm converges.
Proof: Refer to Appendix C-B.
We next find an upper bound on the convergence rate of Algorithm 1.
@@
Consider in addition to the assumptions in Theorem 1, that the step sizes are those of Example 2 with $\upsilon_{1}+\upsilon_{2}<1$. Then, for $\upsilon_{1}=\frac{1}{2}+\frac{\epsilon}{2}$ and $\upsilon_{2}=\frac{1}{6}+\frac{\epsilon}{2}$, we can write
$\begin{split}\frac{\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\mathbb{E}\big{[}\|\nabla F(\theta_{k})\|^{2}\big{]}}{\sum_{k=0}^{K}\alpha_{k}\gamma_{k}}\leq O\left(\frac{1}{K^{\frac{1}{3}-\epsilon}}\right).\\ \end{split}$ (4)
Proof: Refer to Appendix C-C.
@@ IV Experimental results
We classify photos of the two digits “00” and “$1$” from the MNIST dataset using a nonconvex logistic regression model with a regularization parameter of $0.001$. All images are divided equally among the devices and are considered to be preprocessed by being compressed to have dimension $d=10$ using a lossy autoencoder. We run our code on $50$ simulations with different random model initializations testing the accuracy in every iteration against an independent test set. The graphs in Fig. 2 are averaged over all these simulations. For the non-IID data distribution, we first sort the images according to their labels and then divide them among the devices. $\Phi_{k}$ is generated according to Example 1. All channels are generated using the normal distribution with autocovariance $K_{hh}=\frac{1}{2}$. The noise is Gaussian with 00 mean and variance $\sigma_{n}^{2}=\frac{1}{4}$. The step size for FedAvg is taken as $\eta=0.15$. For 1P-ZOFL, $\alpha_{k}=0.5(1+k)^{-0.51}$ and $\gamma_{k}=2.5(1+k)^{-0.18}$. Our algorithm performs consistently well with all the different random variations influencing every simulation in line with our theoretical results. Considering non-IID data distribution seems to be better impacted by our ZO algorithm slightly at the beginning without a major effect on the final result.
In Fig. 3, we test our algorithm with different noise variance $\sigma_{n}^{2}$ and notice that while the algorithm is slowed slightly at the beginning, it does not affect the convergence. When $\sigma_{n}^{2}=2.25$, we change $\alpha_{0}=0.1$ and $\gamma_{0}=0.8$. When $\sigma_{n}^{2}=10.0489$, $\alpha_{0}=0.07$ and $\gamma_{0}=0.3$. We provide a discussion of our algorithm’s performance in relation to the noise variance in Appendix D.
@@ Appendix A On the channel model
$\begin{split}&z\sqrt{2}\cos(2\pi f_{c}t)\\ =&2x\sum_{i}a_{i}\cos(\varphi_{i}(t))\cos^{2}(2\pi f_{c}t)\\ &-2x\sum_{i}a_{i}\sin(\varphi_{i}(t))\sin(2\pi f_{c}t)\cos(2\pi f_{c}t)\\ &+\sqrt{2}w(t)\cos(2\pi f_{c}t)\\ =&x\sum_{i}a_{i}\cos(\varphi_{i}(t))(1+\cos(4\pi f_{c}t))\\ &-x\sum_{i}a_{i}\sin(\varphi_{i}(t))\sin(4\pi f_{c}t)+\sqrt{2}w(t)\cos(2\pi f_{c}t)\\ \end{split}$
@@ B-A Proof of Lemma 1: Biased Estimator
$\begin{split}&\mathbb{E}[g_{k}|\mathcal{H}_{k}]\\ =&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}\Big{(}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(a)}{=}&\mathbb{E}\Big{[}\Phi_{k}\sum_{i=1}^{N}h_{i,k+1}\tilde{F}_{i}\big{(}\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}\big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(b)}{=}&\mathbb{E}\Big{[}\Phi_{k}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{F}_{i}(\theta_{k})\\ &+\gamma_{k}\sum_{i=1}^{N}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}\Phi_{k}^{T}\nabla\tilde{F}_{i}(\theta_{k})\\ &+\gamma_{k}^{2}\sum_{i=1}^{N}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(c)}{=}&\mathbb{E}\Big{[}\Phi_{k}\bigg{(}\frac{\gamma_{k}}{\sigma_{h}^{2}}\sum_{i=1}^{N}h_{i,k+1}h_{i,k}\Phi_{k}^{T}\nabla\tilde{F}_{i}(\theta_{k})\\ &+\gamma_{k}^{2}\sum_{i=1}^{N}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \end{split}$
$\begin{split}=&\frac{\gamma_{k}}{\sigma_{h}^{2}}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}h_{i,k}\Big{|}\mathcal{H}_{k}\Big{]}\mathbb{E}\Big{[}\Phi_{k}\Phi_{k}^{T}\Big{|}\mathcal{H}_{k}\Big{]}\nabla\tilde{F}_{i}(\theta_{k})+\\ &\gamma_{k}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(d)}{=}&\gamma_{k}\beta_{1}\frac{K_{hh}}{\sigma_{h}^{4}}\sum_{i=1}^{N}\nabla F_{i}(\theta_{k})+\\ &\gamma_{k}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(e)}{=}&c_{1}\gamma_{k}(\nabla F(\theta_{k})+b_{k}),\end{split}$ (8)
where $(a)$ is by the definition in (2) and due to Assumption 1, $(b)$ is by Taylor expansion and mean-value theorem and considering $\breve{\theta}_{k}$ between $\theta_{k}$ and $\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}$. $(c)$ is since $\mathbb{E}[h_{i,k+1}]=0$ for the first element and $\mathbb{E}[h_{i,k+1}h_{j,k}]=0$ when $i\neq j$ and the independence of noise for the second element. $(d)$ is due to Assumption 2. In $(e)$, we let $c_{1}=\beta_{1}\frac{K_{hh}}{\sigma_{h}^{4}}$ and the bias $b_{k}=\gamma_{k}\frac{\sigma_{h}^{2}}{\beta_{1}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}\Big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{)}^{2}\Phi_{k}\Phi_{k}^{T}\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\Phi_{k}\Big{|}\mathcal{H}_{k}\Big{]}$.